# Rectified Point Flow: Generic Point Cloud Pose Estimation

### v1

@@ 5 Conclusion
Broader Impact. Rectified Point Flow makes it easier to build reliable 3D alignment and assembly systems directly from raw scans—benefiting robotics, digital manufacturing, AR, and heritage reconstruction. However, the model can still produce incorrect, hallucinated, or nonfunctional assemblies. For safety, further work on assembly verification and error detection will be essential.
Acknowledgments
Tao Sun is supported by the Stanford Graduate Fellowship (SGF). Liyuan Zhu is partially supported by SPIRE Stanford Student Impact Fund Grant. Shuran Song is supported by the NSF Award #2037101. The authors also thank Stanford Marlowe Cluster [62] for providing GPU resources.
@@ Acknowledgments
  • Additional Evaluation (Sec. B): Comparison against category-specific assembly models on PartNet and IKEA-Manual, evaluation on the preservation of rigidity at the part level, and analysis of different generative formulations.
@@
  • Generalization Ability (Sec. 11): Qualitative results on unseen assemblies with same- or cross-category parts to test the model’s generalization ability.
@@ Appendix A Model Details
We adopt a multi-frequency Fourier feature mapping [64], to encode spatial information in both the condition and noised point clouds. For the $j$-th point in the $i$-th part in the condition point cloud, $\mathbf{x}_{i,j}\in\bm{X}$, we construct a 10-dimensional vector, which comprises:
@@ Appendix B Additional Evaluation
We compare against category-specific point-cloud assembly methods in Table 5. All baselines are trained separately for each category, and the category labels are assumed to be known at inference time. RGL-Net [65] additionally assumes a top-to-bottom ordering of the input parts. In contrast, Rectified Point Flow is class-agnostic and performs inference without any class label or part ordering. We evaluated both shape Chamfer Distance (CD) and Part Accuracy (PA) in PartNet-Assembly and IKEA-Manual, following the protocol of Huang et al. [46].
@@
Part-level Rigidity Preservation.
As a dense point map prediction framework, Rectified Point Flow is not explicitly trained to preserve the rigidity of each part. To quantify how well it preserves the rigidity of the parts, we first align each predicted part $\hat{\bm{X}}_{i}(0)$ with the part in assembled state $\bm{X}_{i}(0)$ using the Kabsch algorithm. We then measure two rigidity preservation errors using (1) the Root Mean Square Error (RMSE) over all points and (2) the Overlap Ratio (OR) over all points at varying thresholds $\tau\in\{0.1,0.2,0.5,1,2\}\ \text{cm}$. Specifically, for part $i$ with $M_{i}$ points, we compute
Here, $T^{\prime}_{i}\in\mathrm{SE}(3)$ denotes the optimal rigid transform returned by Kabsch; $\mathbf{x}_{i,j}$ and $\hat{\mathbf{x}}_{i,j}$ denote the $j$-th point on $\bm{X}_{i}(0)$ and on $\hat{\bm{X}}_{i}(0)$, respectively. Because each part is first rigidly aligned to the ground-truth assembled state, these metrics intentionally ignore pose errors, and only measure the shape difference between the predicted and ground truth point parts. To factor in the variations in object size across datasets, we compute the average scale of an object, denoted by $D$, as twice the average distance from the object’s center of gravity to all its points. Then, we define the Relative RMSE as RMSE / $D$, i.e., the RMSE normalized by the average object scale. We report these metrics averaged for all parts in each dataset in Table 6.
Metric Shape Assembly Pairwise Registration
BreakingBad-E TwoByTwo PartNet-A IKEA-M TUD-L ModelNet-40
Object Scale $D$ [cm] – 52.1 107.7 89.0 61.4 40.8 70.0
RMSE [cm] ↓ 0.76 2.46 1.04 0.66 0.16 0.30
Relative RMSE [%] ↓ 1.5 2.3 1.2 1.1 0.4 0.4
OR ($\tau=0.1$ cm) [%] ↑ 52.3 63.8 33.1 46.7 96.9 95.0
OR ($\tau=0.2$ cm) [%] ↑ 61.7 70.8 48.6 57.7 97.1 96.0
OR ($\tau=0.5$ cm) [%] ↑ 74.9 76.8 66.8 69.8 97.4 96.3
OR ($\tau=1$ cm) [%] ↑ 81.4 78.7 77.9 81.0 97.7 96.6
OR ($\tau=2$ cm) [%] ↑ 89.5 81.9 87.4 92.0 98.2 97.1
For the pairwise registration task, Rectified Point Flow demonstrates strong rigidity preservation. On TUD-L, we obtain a Relative RMSE of 0.4% and ORs above 96.9% even at the strictest $\tau=0.1\ \text{cm}$ threshold; on ModelNet-40, we achieve the same Relative RMSE of 0.4% with similar high ORs above 95.0%. Specifically, on TUD-L we record ORs of 96.9% ($\tau=0.1\ \text{cm}$), 97.1% ($\tau=0.2\ \text{cm}$), 97.4% ($\tau=0.5\ \text{cm}$), 97.7% ($\tau=1\ \text{cm}$) and 98.2% ($\tau=2\ \text{cm}$); on ModelNet-40 the corresponding ORs are 95.0%, 96.0%, 96.3%, 96.6% and 97.1%, demonstrating consistently strong rigidity preservation.
In the more challenging shape assembly task, rigidity errors remain low. Across the four datasets, the Relative RMSE ranges from 1.1% to 2.3%. At a strict threshold of $\tau=0.1\,$cm, overlap ratios (ORs) span 33.1 % (PartNet-Assembly) up to 63.8 % (TwoByTwo); By $\tau=1\,$ cm, the ORs exceed 77.9% in the four datasets (77.9%-81.4%), increasing further to 81.9%-92.0% in the more relaxed $\tau=2\,$ cm. The highest Relative RMSE and lower averaged ORs are observed in TwoByTwo, probably due to its limited training samples and lower shape similarity to other datasets, and the fact that TwoByTwo has the largest overall object scale of 107.7 cm among all datasets. In contrast, IKEA-Manual, despite having fewer training samples, benefits from shared priors in furniture objects in joint training, delivering the lowest RMSE and high ORs at all thresholds. These results demonstrate robust rigidity preservation of Rectified Point Flow even in complex shape assembly scenarios.
Note that the subsequent pose recovery stage in Rectified Point Flow further refines part poses via an SVD-based global optimization, which fits optimal poses under noises. In Sec. F, we demonstrate that this SVD-based optimization is crucial to reducing the risk limits of generalization to match the rates achieved by the existing 6-DoF methods. Overall, we empirically confirm that Rectified Point Flow generates point clouds that reliably respect the rigid structure of the conditioning parts.
Ablation on Generative Formulation.
@@
BreakingBad-E TwoByTwo PartNet-A IKEA-M TUD-L ModelNet-40
@@
As an alternative to the generative formulation of Rectified Flow (RF) in our method, we also evaluate a Denoising Diffusion Probabilistic Model (DDPM) [67] using an identical DiT architecture and the pre-trained encoder. In this setup, the forward noising process employs constant variances ($\beta$) that increase linearly from $10^{-4}$ to 0.02 over $T=1000$ timesteps. As shown in Table 7, the RF-based model consistently outperforms the DDPM variant on both shape assembly and pairwise registration tasks, with 35.3% lower rotation error (RE) and 11.63% lower translation error (TE). This result is in line with the findings of GARF [11]. We hypothesize that the straight-line flow in RF reduces the learning difficulty in our tasks. DDPM’s frequency-based generation—which works well for images—may not be as effective as RF for 3D point cloud synthesis in Euclidean space.
@@ Appendix C Randomness in Assembly Generation
To evaluate the diversity of assembly configurations generated by Rectified Point Flow, we sample the Gaussian noise vector $\bm{Z}$ multiple times for the same conditional (unposed) point cloud inputs. At inference time, we set $\bm{X}(1)=\bm{Z}$ and run the model to obtain prediction $\hat{\bm{X}}(0)$. In Figure 9, each row corresponds to a single final assembly: the first column shows the ground-truth assembly, and the next four columns display outputs produced by four different Gaussian noises. All generated assemblies preserve the part structure, yet exhibit meaningful variations in the parts’ placement and orientation, and overall geometry of the object. As expected, the model produces diverse configurations for symmetric or interchangeable parts, such as the armrests and the chair base. This shows that Rectified Point Flow effectively captures a diverse conditional distribution of valid assemblies.
@@ Appendix D Generalization Ability
Figure 11: Generalization to Unseen Assemblies Within the Same Category: We select parts from two objects of the same category in the PartNet-Assembly test set. Parts from Object 1 are shown in blue, and parts from Object 2 in red; unselected parts are shown in gray. The results demonstrate that the model comprehends the underlying geometric structure of the category and can re-target parts to construct the final shape of the same category.
@@
As shown in Figure 11, parts selected from Object 1 are rendered in blue and those from Object 2 in red. Our model then synthesizes novel assemblies that blend and reconfigure these parts in a coherent and category-consistent manner. For example, in the chair category (first two rows), the model successfully retains a functional and plausible seat-back-leg structure while creatively mixing parts. In the lamp category (third row), even though the base and shade style differ significantly between objects, generated results exhibit sensible combinations that maintain structural integrity. Similarly, in the table category (last row), our method combines parts from a flat-top table and lattice-style base to produce hybrid yet coherent table designs.
@@
Figure 12 highlights Rectified Point Flow’s ability to generalize to unseen part combinations across categories. This is a particularly challenging test, since such part combination may not even be possible to be assembled into a meaningful object. Nevertheless, our method still demonstrates a certain degree of generalization. We show two input objects from different categories, for example, a monitor and a chair, a chair and a lamp, or a wall sconce and a spray bottle. The results on the right demonstrate that our model can reconfigure these parts into plausible new assemblies, preserving geometric coherence. This suggests that the model has learned a strong understanding of part relationship, allowing it to reason about compositionality even across category boundaries.
@@ Appendix E Proof of Theorem 1
Lemma 1 ($\mathcal{G}$-invariance of the flow distribution).
@@
  • Rotational symmetry: If $R\in G_{i}$, then $R\bm{X}_{i}(0)=\bm{X}_{i}(0)$ point-wise, so $p(\mathbf{x}(0))=p(R\mathbf{x}(0))$.
@@ Appendix F Generalization Bounds
While the Rectified Point Flow predicts a much higher-dimensional space ($3M_{i}$ coordinates per part), we find that its Rademacher complexity scales exactly the same rate as the 6-DoF methods, $O(1/\!\sqrt{m})$, where $m$ is the number of samples in the training set.
@@
^$\hat{\mathcal{R}}_{D}(h)=\frac{1}{m}\sum_{k=1}^{m}\ell\bigl{(}h(C^{(k)}),R^{\star(k)},\bm{t}^{\star(k)}\bigr{)}.$
@@
Let $\bm{X}^{\star}\!\in\!\mathbb{R}^{3N}$ be the centralized ground-truth point set of a single part, and denote $\sigma_{\min}=\sigma_{\min}((\bm{X}^{\star})^{\top}\bm{X}^{\star})$. If $\|\hat{\bm{X}}(0)-\bm{X}^{\star}\|_{F}\leq\varepsilon$, the optimal Procrustes solution $(\hat{R},\hat{\bm{t}})=P(\hat{\bm{X}}(0))$ satisfies
@@
This property directly follows from Davis–Kahan perturbation bounds [68] for the Top-$3$ singular vectors. Crucially, $\sigma_{\min}=\Omega(N)$ [^2] Here, $\Omega(\cdot)$ denote the asymptotic rate, instead of part index set. for well-spread point clouds, so $\mathrm{Pr}$ is a $\tfrac{1}{\sqrt{N}}$-Lipschitz map.
Let $\mathfrak{R}_{m}(\mathcal{H})$ denote the empirical Rademacher complexity on $S$. Because composition with a $L$-Lipschitz map contracts Rademacher complexity
@@
$\mathcal{R}\bigl{(}P\!\circ\!\hat{f}\bigr{)}$$\;\leq\;\hat{\mathcal{R}}_{D}\bigl{(}P\!\circ\!\hat{f}\bigr{)}+2\,\mathfrak{R}_{m}(P\!\circ\!\mathcal{F})+3\,\sqrt{\frac{\log(2/\delta)}{2m}},$ (Flow)
$\mathcal{R}(\hat{g})$$\;\leq\;\hat{\mathcal{R}}_{D}(\hat{g})+2\,\mathfrak{R}_{m}(\mathcal{G})+3\,\sqrt{\frac{\log(2/\delta)}{2m}},$ (6DoF)
where $\hat{f}\!\in\!\mathcal{F}$ and $\hat{g}\!\in\!\mathcal{G}$ are the empirical-risk minimizers on $S$.
In conclusion, while Rectified Point Flow predicts a much higher-dimensional space, the contraction of the SVD stage cancels this apparent over-parameterization, producing a complexity term that scales at the same rate of $O(1/\!\sqrt{m})$ as the 6-DoF baseline; (Flow)–(6DoF).
As a result, our method enjoys at least same generalization risk guarantees despite operating in an over-parameterized prediction space, while retaining the $\mathcal{G}$-invariance benefits proven in Sec. E.

### v2

@@ 5 Conclusion
Broader Impact. Rectified Point Flow makes it easier to build reliable 3D alignment and assembly systems straight from raw scans, benefiting robotics, digital manufacturing, AR, and heritage reconstruction. Given its performance and speed, it reduces the barrier for applying 3D part reasoning in resource-constrained settings. However, the model can still produce incorrect, hallucinated, or nonfunctional assemblies. For safety, further work on assembly verification and assembly error recovery will be essential.
Acknowledgments and Disclosure of Funding
Tao Sun is supported by the Stanford Graduate Fellowship (SGF). Liyuan Zhu is partially supported by SPIRE Stanford Student Impact Fund Grant. Shuran Song is supported by the NSF Award #2037101. We appreciate the authors of GARF [11] for providing the Fractura dataset. We also thank Stanford Marlowe Cluster [64] for providing GPU resources.
@@
  • Additional Evaluation (Sec. B):
    • Runtime analysis.
    • Evaluation on the preservation of rigidity at the part level.
    • Comparison against category-specific assembly models on PartNet and IKEA-Manual.
    • Analysis of different generative formulations,
    • Evaluation of the anchor-free version of our model.
@@
  • Generalization Ability (Sec. D): Qualitative results on unseen assemblies with same- or cross-category parts to test the model’s generalization ability.
@@ Appendix A Model Details
We adopt a multi‐frequency Fourier feature mapping [66], to encode spatial information in both the condition and noised point clouds. For the $j$-th point in the $i$-th part in the condition point cloud, $\mathbf{x}_{i,j}\in\bm{X}$, we construct a 10-dimensional vector, which comprises:
@@ Appendix B Additional Evaluation
Runtime Analysis.
The number of sampling steps of the flow model is an important factor that affects both accuracy and runtime. In Tab. 7, we vary the sampling steps and report the Part Accuracy, Chamfer Distance, and the runtime per sample in PartNet-Assembly, measured on a single RTX 4090 GPU. Increasing steps consistently improves accuracy and geometric precision, with diminishing returns beyond 20 steps. We use $K=20$ sampling steps for all evaluations in the paper. In this setting, our model achieves 4.3 samples per second, making it practical for robotic assembly and localization tasks that require frequent online inference.
Table 7: Trade-offs of Sampling Steps between Accuracy and Runtime. Metrics and runtime evaluated on PartNet-Assembly dataset with a single NVIDIA RTX 4090 GPU.
Sampling steps 1 2 5 10 20 50
Part Accuracy [%] $\uparrow$ 25.2 38.1 46.7 52.1 53.9 54.6
Chamfer Distance [cm] $\downarrow$ 3.23 1.70 0.90 0.75 0.73 0.71
Runtime / sample [s] $\downarrow$ 0.072 0.081 0.108 0.148 0.232 0.483
Part-level Rigidity Preservation.
As a dense point map prediction framework, Rectified Point Flow is not explicitly trained to preserve the rigidity of each part. To quantify how well it preserves the rigidity of the parts, we first align each predicted part $\hat{\bm{X}}_{i}(0)$ with the part in assembled state $\bm{X}_{i}(0)$ using the Kabsch algorithm.
We then measure two rigidity preservation errors using (1) the Root Mean Square Error (RMSE) over all points and (2) the Overlap Ratio (OR) over all points at varying thresholds $\tau\in\{0.1,0.2,0.5,1,2\}\ \text{cm}$. Specifically, for part $i$ with $M_{i}$ points, we compute
Here, $T^{\prime}_{i}\in\mathrm{SE}(3)$ denotes the optimal rigid transform returned by Kabsch; $\mathbf{x}_{i,j}$ and $\hat{\mathbf{x}}_{i,j}$ denote the $j$-th point on $\bm{X}_{i}(0)$ and on $\hat{\bm{X}}_{i}(0)$, respectively. Because each part is first rigidly aligned to the ground-truth assembled state, these metrics intentionally ignore pose errors, and only measure the shape difference between the predicted and ground truth point parts. To factor in the variations in object size across datasets, we compute the average scale of an object, denoted by $D$, as twice the average distance from the object’s center of gravity to all its points. Then, we define the Relative RMSE as RMSE / $D$, i.e., the RMSE normalized by the average object scale. We report these metrics averaged for all parts in each dataset in Tab. 8.
Metric Shape Assembly Pairwise Registration
BreakingBad-E TwoByTwo PartNet-A IKEA-M TUD‐L ModelNet‐40
Object Scale $D$ [cm] – 52.1 107.7 89.0 61.4 40.8 70.0
RMSE [cm] ↓ 0.76 2.46 1.04 0.66 0.16 0.30
Relative RMSE [%] ↓ 1.5 2.3 1.2 1.1 0.4 0.4
OR ($\tau=0.1$ cm) [%] ↑ 52.3 63.8 33.1 46.7 96.9 95.0
OR ($\tau=0.2$ cm) [%] ↑ 61.7 70.8 48.6 57.7 97.1 96.0
OR ($\tau=0.5$ cm) [%] ↑ 74.9 76.8 66.8 69.8 97.4 96.3
OR ($\tau=1$ cm) [%] ↑ 81.4 78.7 77.9 81.0 97.7 96.6
OR ($\tau=2$ cm) [%] ↑ 89.5 81.9 87.4 92.0 98.2 97.1
For the pairwise registration task, Rectified Point Flow demonstrates strong rigidity preservation. On TUD-L, we obtain a Relative RMSE of 0.4% and ORs above 96.9% even at the strictest $\tau=0.1\ \text{cm}$ threshold; on ModelNet-40, we achieve the same Relative RMSE of 0.4% with similar high ORs above 95.0%. Specifically, on TUD‑L we record ORs of 96.9% ($\tau=0.1\ \text{cm}$), 97.1% ($\tau=0.2\ \text{cm}$), 97.4% ($\tau=0.5\ \text{cm}$), 97.7% ($\tau=1\ \text{cm}$) and 98.2% ($\tau=2\ \text{cm}$); on ModelNet-40 the corresponding ORs are 95.0%, 96.0%, 96.3%, 96.6% and 97.1%, demonstrating consistently strong rigidity preservation.
In the more challenging shape assembly task, rigidity errors remain low. Across the four datasets, the Relative RMSE ranges from 1.1% to 2.3%. At a strict threshold of $\tau=0.1\,$cm, overlap ratios (ORs) span 33.1 % (PartNet‑Assembly) up to 63.8 % (TwoByTwo); By $\tau=1\,$ cm, the ORs exceed 77.9% in the four datasets (77.9%-81.4%), increasing further to 81.9%-92.0% in the more relaxed $\tau=2\,$ cm. The highest Relative RMSE and lower averaged ORs are observed in TwoByTwo, probably due to its limited training samples and lower shape similarity to other datasets, and the fact that TwoByTwo has the largest overall object scale of 107.7 cm among all datasets. In contrast, IKEA-Manual, despite having fewer training samples, benefits from shared priors in furniture objects in joint training, delivering the lowest RMSE and high ORs at all thresholds. These results demonstrate robust rigidity preservation of Rectified Point Flow even in complex shape assembly scenarios.
Furthermore, please note that the subsequent pose recovery stage in Rectified Point Flow further refines part poses via an SVD-based global optimization, which fits optimal poses under noises. Overall, we empirically confirm that Rectified Point Flow generates point clouds that reliably respect the rigid structure of the conditioning parts.
@@
We compare against category-specific point‐cloud assembly methods in Tab. 9. All baselines are trained separately for each category, and the category labels are assumed to be known at inference time. RGL-Net [67] additionally assumes a top-to-bottom ordering of the input parts. In contrast, Rectified Point Flow is class-agnostic and performs inference without any class label or part ordering. We evaluated both shape Chamfer Distance (CD) and Part Accuracy (PA) in PartNet-Assembly and IKEA-Manual, following the protocol of Huang et al[47].
@@
BreakingBad-E TwoByTwo PartNet-A IKEA-M TUD‐L ModelNet‐40
@@
Ablation on Generative Formulation.
As an alternative to the generative formulation of Rectified Flow (RF) in our method, we also evaluate a Denoising Diffusion Probabilistic Model (DDPM) [69] using an identical DiT architecture and the pre-trained encoder. In this setup, the forward noising process employs constant variances ($\beta$) that increase linearly from $10^{-4}$ to 0.02 over $T=1000$ timesteps. As shown in Tab. 10, the RF-based model consistently outperforms the DDPM variant on both shape assembly and pairwise registration tasks, with 35.3% lower rotation error (RE) and 11.63% lower translation error (TE). This result is in line with the findings of GARF [11]. We hypothesize that the straight-line flow in RF reduces the learning difficulty in our tasks. DDPM’s frequency‑based generation—which works well for images—may not be as effective as RF for 3D point cloud synthesis in Euclidean space.
Anchor-free Models.
In the anchor-free setting, we do not fix the anchor’s pose. Instead, during training, we treat the anchor exactly like any other part in the conditioning: its point cloud is centered to its own CoM frame and then randomly rotated. The model, therefore, never receives the true anchor pose from condition. The flow target is defined in the anchor frame: we put the completed assembly point cloud in the anchor’s frame and then re-center the whole assembled point cloud. At inference, we first estimate the rigid transform between the predicted anchor and the ground-truth anchor point cloud via ICP, and then apply this transform to all predicted parts. Metrics (RE, TE, and Part Acc.) are computed after this anchor alignment.
Across shape assembly datasets (Tab. 11) and pairwise registration (Tab. 12), anchor-free (average) underperforms anchor-fixed as expected due to (i) error propagation from anchor misalignment and (ii) ambiguity induced by symmetric anchor parts. However, despite the lower absolute performance, our anchor-free model preserves the same relative ordering among baselines on all shape-assembly datasets; the only exception is GARF, which is evaluated in the anchor-fixed protocol. In particular, our model significantly improves anchor-free SOTA on BreakingBad-Everyday’s Part Accuracy from 76.2% (PuzzleFussion++ [25]) to 90.2%. For pairwise registration, our model is also achieving the lowest TE on ModelNet and the highest Recall@5 on TUD-L among baselines, while remain competitive on other metrics.
Furthermore, we observe that the anchor-free (best-of-3) evaluation, which selects the best prediction out of 3 different random seeds per sample, largely narrows the gap to anchor-fixed. This indicates that a small stochastic budget may find a more reliable anchor-free prediction. The effect is most pronounced on datasets with higher part count and weaker geometric cues (e.g., PartNet-Assembly and IKEA-Manual), where anchor ambiguity is stronger and small anchor errors propagate severely.
Table 11: Evaluation of the Anchor-free Model on Shape Assembly Datasets.
Dataset Anchor-fixed Anchor-free (Average) Anchor-free (Best-of-3)
RE $\downarrow$ TE $\downarrow$ PA $\uparrow$ RE $\downarrow$ TE $\downarrow$ PA $\uparrow$ RE $\downarrow$ TE $\downarrow$ PA $\uparrow$
[deg] [cm] [%] [deg] [cm] [%] [deg] [cm] [%]
BreakingBad-E [12] 7.4 2.0 93.5 17.4 8.0 90.2 13.0 5.9 92.2
TwoByTwo [16] 13.2 3.0 15.2 24.2 9.0 16.7
PartNet-Assembly [15] 21.8 14.8 53.9 47.3 40.5 45.3 38.2 32.9 54.3
IKEA-Manual [17] 20.7 24.7 33.2 54.7 51.5 19.5 39.0 40.5 28.6
Table 12: Evaluation of the Anchor-free Model on Pairwise Registration Datasets.
Setting ModelNet-40 [19] TUD-L [18]
RE $\downarrow$ TE $\downarrow$ Recall @ 5 $\uparrow$ Recall @ 1cm $\uparrow$
[deg] [unit] [%] [%]
Anchor-fixed 0.93 0.002 97.7 99.0
Anchor-free (Average) 2.05 0.012 96.6 96.3
Anchor-free (Best-of-3) 1.18 0.007 97.3 97.3
@@ Appendix C Randomness in Assembly Generation
To evaluate the diversity of assembly configurations generated by Rectified Point Flow, we sample the Gaussian noise vector $\bm{Z}$ multiple times for the same conditional (unposed) point cloud inputs. At inference time, we set $\bm{X}(1)=\bm{Z}$ and run the model to obtain prediction $\hat{\bm{X}}(0)$. In Figure 9, each row corresponds to a single final assembly: the first column shows the ground‐truth assembly, and the next four columns display outputs produced by four different Gaussian noises. All generated assemblies preserve the part structure, yet exhibit meaningful variations in the parts’ placement and orientation, and overall geometry of the object. As expected, the model produces diverse configurations for symmetric or interchangeable parts, such as the armrests and the chair base. This shows that Rectified Point Flow effectively captures a diverse conditional distribution of valid assemblies.
Linear Interpolation in Noise Space.
@@ Appendix D Generalization Ability
Figure 11: Generalization to Unseen Assemblies Within the Same Category: We select parts from two objects of the same category in the PartNet-Assembly test set. Parts from Object 1 are shown in blue, and parts from Object 2 in red; unselected parts are shown in gray. The results demonstrate that the model comprehends the underlying geometric structure of the category and can re-target parts to construct the final shape.
@@
As shown in Fig. 11, parts selected from Object 1 are rendered in blue and those from Object 2 in red. Our model then synthesizes novel assemblies that blend and reconfigure these parts in a coherent and category-consistent manner. For example, in the chair category (first two rows), the model successfully retains a functional and plausible seat-back-leg structure while creatively mixing parts. In the lamp category (third row), even though the base and shade style differ significantly between objects, generated results exhibit sensible combinations that maintain structural integrity. Similarly, in the table category (last row), our method combines parts from a flat-top table and lattice-style base to produce hybrid yet coherent table designs.
@@
Fig. 12 highlights Rectified Point Flow’s ability to generalize to unseen part combinations across categories. This is a particularly challenging test, since such part combination may not even be possible to be assembled into a meaningful object. Nevertheless, our method still demonstrates a certain degree of generalization. We show two input objects from different categories, for example, a monitor and a chair, a chair and a lamp, or a wall sconce and a spray bottle. The results on the right demonstrate that our model can reconfigure these parts into plausible new assemblies, preserving geometric coherence. This suggests that the model has learned a strong understanding of part relationship, allowing it to reason about compositionality even across category boundaries.
@@ Appendix E Proof of Theorem 1
Lemma 1 ($\mathcal{G}$‑invariance of the flow distribution).
@@
  • Rotational symmetry: If $R\in G_{i}$, then $R\bm{X}_{i}(0)=\bm{X}_{i}(0)$ point‑wise, so $p(\mathbf{x}(0))=p(R\mathbf{x}(0))$.
@@ Appendix F Generalization Bounds
While the Rectified Point Flow predicts a much higher‑dimensional space ($3M_{i}$ coordinates per part), we find that its Rademacher complexity scales exactly the same rate as the 6‑DoF methods, $O(1/\!\sqrt{m})$, where $m$ is the number of samples in the training set.
@@
$\hat{\mathcal{R}}_{D}(h)=\frac{1}{m}\sum_{k=1}^{m}\ell\bigl(h(C^{(k)}),R^{\star(k)},\bm{t}^{\star(k)}\bigr).$
@@
Let $\bm{X}^{\star}\!\in\!\mathbb{R}^{3N}$ be the centralized ground‑truth point set of a single part, and denote $\sigma_{\min}=\sigma_{\min}((\bm{X}^{\star})^{\top}\bm{X}^{\star})$. If $\|\hat{\bm{X}}(0)-\bm{X}^{\star}\|_{F}\leq\varepsilon$, the optimal Procrustes solution $(\hat{R},\hat{\bm{t}})=P(\hat{\bm{X}}(0))$ satisfies
@@
This property directly follows from Davis–Kahan perturbation bounds [70] for the Top‑$3$ singular vectors. Crucially, $\sigma_{\min}=\Omega(N)$ [^2] Here, $\Omega(\cdot)$ denote the asymptotic rate, instead of part index set. for well‑spread point clouds, so $\mathrm{Pr}$ is a $\tfrac{1}{\sqrt{N}}$‑Lipschitz map.
Let $\mathfrak{R}_{m}(\mathcal{H})$ denote the empirical Rademacher complexity on $S$. Because composition with a LL‑Lipschitz map contracts Rademacher complexity
@@
[…]
where $\hat{f}\!\in\!\mathcal{F}$ and $\hat{g}\!\in\!\mathcal{G}$ are the empirical‑risk minimizers on $S$.
In conclusion, while Rectified Point Flow predicts a much higher‑dimensional space, the contraction of the SVD stage cancels this apparent over-parameterization, producing a complexity term that scales at the same rate of $O(1/\!\sqrt{m})$ as the 6‑DoF baseline; (Flow)–(6DoF).
As a result, our method enjoys at least same generalization risk guarantees despite operating in an over‑parameterized prediction space, while retaining the $\mathcal{G}$-invariance benefits proven in Sec. E.